OpenStudy (anonymous):

Letf(x)=-4x+7 and g(x)=2x-6. Find (f (open O) g) (1)

OpenStudy (anonymous):

ok now we got it. first compute \[g(1)=2\times 1-6=2-6=-4\]

OpenStudy (anonymous):

then compute \[f(-4)=-4\times (-4)+7=16+7=23\]

OpenStudy (anonymous):

or compute \[f(g(x))=f(2x-6)=-4(2x-6)+7=-8x+24+7=-8x+31\]

OpenStudy (anonymous):

then \[f(g(1))=-8\times 1+31=-8+31=23\]

OpenStudy (anonymous):

your choice

OpenStudy (anonymous):

23?

OpenStudy (anonymous):

yes 23